18.01 Syllabus, Spring Term 1997

Text: Calculus with Analytic Geometry, 2nd Edition, by George F. Simmons, McGraw-Hill, and the 18.01 Supplementary Notes and Problems from Jerison’s course, Fall 1996.

Lecture: MWF 1 in 2–142 Fred Diamond, 2–278, 253–4996, fdiamond@math.mit.edu

Office Hours: To be announced in class.

Recitation: TR 1 in 2–135 Jim Rath, 4–174, 253–2634, ratjamm@mit.edu

Office Hours: M3-5 and F9-12 or by appt.

Tutoring: Tutoring is available in Room 2–102. It is open M-R 3-5 and 7:30-9:30. Tutoring is also available from the Tutorial Services Room (12–124, 253–8406).

Homework: Homework will be collected about once a week. Assignments will be handed out in class. You may collaborate on homework, but when you write out your solutions, you must work alone and without consulting solutions written by anyone else.

Exams: There will be 5 in-class exams, scheduled for Feb. 21, March 14, April 4, April 25 and May 9. If it you are unable to take an exam at the scheduled time, you must contact me as soon as possible to make other arrangements. If you fail an in-class exam, you can try again by arranging to take a “tutored exam” during the following week. The tutored exam is then worth at most the minimum passing grade.

Grading: Your grade will be determined by: in-class tests, 500 points; homework, 250 points; final exam, 250 points.